I need help solving a linear system problem using linear comblinations. The problem is 7x-4y=-3 and 2x+5y=-7. I got (1,1), but when I use them they don't work right. First I multiplied 5(7x-4y)=5(-3), 35y-20y=-15, 43y=43, y-1.
The next part I multiplied 2(2x+5y)=2(-7), 14x-8y=-6, -14x-35y=49, 43y=43, y=1.
I'm not sure where I need to go with all of this. Please help me.

Found 2 solutions by nyc_function, richwmiller:
Answer by nyc_function(2741)

7x-4y=-3...Equation A
2x+5y=-7...Equation B
7x - 4y = -3
-4y = -7x - 3
y = (7/4)x + 3/4
Plug this into Equation B to find x.
2x + 5((7/4)x + 3/4) = -7
2x + (35/4)x + 15/4 = -7
2x(4) + 4[(35/4)x + 15/4] = -7(4)
8x + 35x + 3 = -28
43x = -28 - 3
43x = -31
x = -31/43
==================================
Now, plug this value for x into either equations above to find y.
I will use Equation A but you can also use B.
7(-31/43) - 4y = -3
-217/43 - 4y = -3
-4y = -3 + 217/43
y = (-3 + 217/43)/-4
y = -22/43
SOLUTION POINT: (-31/43, -22/43)....This is the answer.
The point (1,1) is not the answer.



Answer by richwmiller(17219)

the other tutor showed you the right wayn tosolve it.
Let's see where you went wrong.
5(7x-4y)=5(-3)
35y-20y=-15,
This is where you went wrong. It should be
35x-20y=-15
You changed x to y
Now for the other equation
2(2x+5y)=2(-7)
should be
4x+10y-14
not sure where you got this
14x-8y=-6
or this
-14x-35y=49
I also don't see what you were trying to do by multiplying the first by 5 and the second by 2. I can see multiplying the first by 5 and the second by 4 to remove the y term -20y and +20y